OpenStudy (anonymous):

Approximate the volume of a spherical shell where the inner radius is 4 inches and whose thickness is 1/16 inches using DIFFERENTIALS.

OpenStudy (anonymous):

hi. do you still need help with this problem?

OpenStudy (anonymous):

yes. will you help me?

OpenStudy (anonymous):

i will do what i can, yes

OpenStudy (anonymous):

so let's start off with whether or not we can find some applicable formulas here.

OpenStudy (anonymous):

your heaven sent thank you..:)

OpenStudy (anonymous):

do you have the answer to this thing?

OpenStudy (anonymous):

none..

OpenStudy (anonymous):

k. Let's start with what we know. Like maybe the volume of a sphere.

OpenStudy (anonymous):

Do you remember the formula for the volume of a sphere? dont worry, you are gonna get through this problem. I promise

OpenStudy (anonymous):

(pi)(d)^3/6

OpenStudy (anonymous):

or 4/3 (pi) R^3

OpenStudy (anonymous):

yes... that is what we seek..

OpenStudy (anonymous):

now what is the variable that affects the volume? Because that is the variable that will drive the differentiation process.

OpenStudy (anonymous):

Radius?

OpenStudy (anonymous):

absolutely.. good. Now how do we get the differential of this little puppy?

OpenStudy (anonymous):

solve for V' ?

OpenStudy (anonymous):

what is it we are approximating?

OpenStudy (anonymous):

wait. i didnt see the prime

OpenStudy (anonymous):

the volume of a spherical shell

OpenStudy (anonymous):

yes. we want V'

OpenStudy (anonymous):

so how does one go about getting V'?

OpenStudy (anonymous):

in reality, it is dV/dr we need.

OpenStudy (anonymous):

\[\frac{ dv }{ dr }=(\frac{ 4 }{ 3 })(\pi)(3R)\] not sure..is it right?

OpenStudy (anonymous):

you are very close. check the exponent on the R term

OpenStudy (anonymous):

3R^2 :)

OpenStudy (anonymous):

yes. so the whole differential looks like.... dV=

OpenStudy (anonymous):

remember that dV=(dV/dr)dr

OpenStudy (anonymous):

dv = (4/3)(pi)(3R^2)dr ?

OpenStudy (anonymous):

perfect!

OpenStudy (anonymous):

notice how the 3's cancel

OpenStudy (anonymous):

for somebody who desperately thought you could nt do this problem, you are doing remarkably well :)

OpenStudy (anonymous):

dv = 4 pi R^2 dr

OpenStudy (anonymous):

yes. now plug in your given values and u are done. excellent work

OpenStudy (anonymous):

notice how you can approximate the change in volume by taking the product of the surface area of the sphere and dr.

OpenStudy (anonymous):

where do i substitute 4 inches to? 1/16 inches?

OpenStudy (anonymous):

r=4 and dr=1/16

OpenStudy (anonymous):

dv = 12.5663... ?

OpenStudy (anonymous):

this approximation is good because the dr or differential of r is very small ( 1/16)

OpenStudy (anonymous):

ummm... i dont have a calc with me. but that looks pretty close to 4pi, which is the answer i got

OpenStudy (anonymous):

thank you (1000000000x)!!!

OpenStudy (anonymous):

it was such a pleasure

OpenStudy (anonymous):

wait one last question...

OpenStudy (anonymous):

sure. shoot.

OpenStudy (anonymous):

find the value of cos 43 using differentials...

OpenStudy (anonymous):

lol. i think i saw that problem an hour or two ago.

OpenStudy (anonymous):

yes..but no one helped me :(

OpenStudy (anonymous):

i am so sorry. i saw someone looking so i decided to move on.

OpenStudy (anonymous):

ok. the drill is the same....except... ummm... slightly different. and no, that is not a DIFFERENTial joke.

OpenStudy (anonymous):

so where do you think you might start? what is the reference angle you would use in this case? 43 degrees is very close to .....

OpenStudy (anonymous):

45 degrees

OpenStudy (anonymous):

good. that is exactly correct

OpenStudy (anonymous):

what function are we differentiating?

OpenStudy (anonymous):

cos

OpenStudy (anonymous):

y=....

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

now i could be mistaken, but i think we need to do this in radians

OpenStudy (anonymous):

so? what's the next step?

OpenStudy (anonymous):

find the formula for the differential given that y=cos x

OpenStudy (anonymous):

dy/dx = -sin x ?

OpenStudy (anonymous):

perfect. and so dy=?

OpenStudy (anonymous):

-sin x dx

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

x=? dx?

OpenStudy (anonymous):

no, dy=-sin x dx

OpenStudy (anonymous):

i mean what is x=? dx=?

OpenStudy (anonymous):

x=45=pi/4 and dx=-2=-pi/90

OpenStudy (anonymous):

why -pi/90?

OpenStudy (anonymous):

cuz that is 2 degrees in radians. all trig function are calculated in radians in calc. cant remember for the life of me why tho. sorry about that

OpenStudy (anonymous):

(-2)xpi/180=-pi/90. the pi/180 converts degrees to radians

OpenStudy (anonymous):

remember that dx=delta x=43-45=-2

OpenStudy (anonymous):

ahh..gets

OpenStudy (anonymous):

if you are not comfy cozy with this, please feel free to ask me questions

OpenStudy (anonymous):

is dx= change in x?

OpenStudy (anonymous):

yes, by the definition of a differential. the differential of y, dy, equals y'x dx, where dx=the change in x

OpenStudy (anonymous):

oops.. the differential of y still equals y'x dx, not y'x dx.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

y'dx

OpenStudy (anonymous):

there we go

OpenStudy (anonymous):

oh okay...thanks a lot again!! i hope you'll never get tired of helping me.. see ya next time??

OpenStudy (anonymous):

i hope so. it has been a pleasure. you take care unk. lol. bye